I have a parabolic shape that is shown as: y = 2x^2
I also need to graph an ellipse: (y - 2)^2/1 + x^2/4 = 1
Then, I need to decide if the parabola fits in the hole.
: I have a parabolic shape that is shown as: y = 2x^2
I also need to graph an ellipse: (y - 2)^2/1 + x^2/4 = 1
Then, I need to decide if the parabola fits in the hole.

Answer by Nate(3495)

graph(300,300,-5,5,-5,5,2x^2,sqrt(1 - x^2/4) + 2,-sqrt(1 - x^2/4) + 2)
I guess it does fit if you mean 3D. But it is just plain in 2D.
Question 58576: I have a parabolic shape that is shown as: y = 2x^2
I also need to graph an ellipse: (y - 2)^2/1 + x^2/4 = 1
Then, I need to decide if the parabola fits in the hole.
: I have a parabolic shape that is shown as: y = 2x^2
I also need to graph an ellipse: (y - 2)^2/1 + x^2/4 = 1
Then, I need to decide if the parabola fits in the hole.

Answer by stanbon(19670)

I have a parabolic shape that is shown as: y = 2x^2
I also need to graph an ellipse: (y - 2)^2/1 + x^2/4 = 1
Then, I need to decide if the parabola fits in the hole.
----------------
You need to carefully sketch the parabolic shape and the
information about the ellipse on a graph. Then make a
determination regarding the answer to your question.
-----------------
Your parabolic shape has a vertex at (0,0) but you haven't
said what it's domain is---that would determine how far up
or down or to he right or to the left it goes.
--------
The ellipse has its center at (0,2).
------------
When x=0, (y-2)^2=1 ; y-2=+-1; y=3 or y=1
So the ellipse intersects the y-axis at (0,1) and (0,3)
---------
When y=2, x^2=4; x=+-2
So the ellipse passes through (2,2) and (-2,2)
When y=0, 4+x^2/4=1; x^2=-12;
So the ellipse does not touch the x-axis.
----------
Hopefully from all this you can figure out the
answer to your question.
Cheers,
Stan H.